OpenStudy (anonymous):

If you have a device that has an energy efficiency of 0.4 and you put in 100 units of energy, how many units of useful energy will you get out of it?

OpenStudy (kropot72):

\[\large Energy\ efficiency=\frac{energy\ output}{energy\ input}\] Therefore we can write: energy output = (energy input) * (efficiency) = 100 * 0.4 = you can calculate